• Main INDEX
  • Monthly INDEX
  • PREV
  • NEXT

    User name J. Singh

    Log entry time 08:33:06 on August22,2003

    Entry number 112592

    Followups:

    keyword=owl shift summary

    owl shift run summary is forthcoming.


    0000:0155 Finishing/Backing out of Moller Measurement.
    0155:0202 Waiting for beam (first BCM trip)
    0202:0220 Harp scans x 4 = 166x144 microns
    0220:0305 Raster spot check
    0305:0329 BCM trips, moving to carbon multi
    0329:0337 accept, 0%, carbon multi, raster ON
    0337:0339 raster = OFF
    0339:0347 accept, 0%, carbon multi, raster OFF, y_target spectra
    0347:0353 move to empty target
    0353:0359 accept, 0%, empty target, raster OFF
    0359:0402 raster = ON
    0402:0408 accept, 0%, empty target, raster ON
    0408:0416 move to ref cell
    0416:0421 accept, 0%, ref cell = empty, raster ON
    0421:0426 filling w/ N2, beam on = 140 psig, beam off = 138 psig
    0426:0431 accept, 0%, ref cell = N2, raster ON
    0431:0455 changing to -3.6%, moving to carbon multi, y_target spectra
    0455:0502 accept, -3.6%, carbon multi, raster OFF
    0502:0508 move to empty target
    0508:0514 accept, -3.6%, empty target, raster OFF
    0514:0536 changing to -7.2%, moving to carbon multi
    0536:0543 accept, -7.2%, carbon multi, raster OFF
    0543:0549 move to empty target
    0549:0555 accept, -7.2%, empty target, raster OFF
    0555:0800 changing to kin 12.10, controlled access, carbon foil damage